# SmallToLarge (S2L): Scalable Data Selection for Fine-tuning Large Language Models by Summarizing Training Trajectories of Small Models

### v1

@@ 5 Experiments
In this section, we present the comprehensive experiments conducted to evaluate the efficacy of the proposed data selection method, SmallToLarge (S2L), across two challenging domains (mathematical reasoning and clinical text summarization) and various model sizes (from 410M to 2.8B) and architectures (Pythia, Phi). Additionally, we incorporate ablation studies to further understand S2L, including the effect of length and timing of the trajectory. We fine-tune all the models with the Huggingface transformers library (Wolf et al., 2020) with Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) on 4 48G NVIDIA RTX A6000.
@@ 5.1 Baselines
We systematically compare S2L against a comprehensive set of open-sourced data selection methods. These methods are categorized based on the type of representation they use and selected as the most representative or best-performing methods as identified in prior work. These include: (1) Random Sampling; (2) Least Confidence selection (Bhatt et al., 2024); (3) selection based on Middle Perplexity (Marion et al., 2023); (4) High Learnability, determined by the loss decrease before and after full fine-tuning (Zhou et al., 2023b); and (5) Facility Locations selection based on hidden states (Bhatt et al., 2024). Additionally, we incorporate two online selection techniques: (6) DiverseEvol (Wu et al., 2023b), which progressively adds examples with distinct embeddings, and the (7) Confidence Curriculum proposed by Varshney et al., which uses the confidence score averaged over the past few epochs to select examples with decreasing confidence during the training while mixing in a certain fraction of higher confidence examples selected in the previous rounds.
Given that the optimal reference model may vary for each one-shot selection method, we ensure a fair comparison by adopting the approach used in (Marion et al., 2023), which runs each method with both the fully fine-tuned target model and an early fine-tuning checkpoint as the reference model. We report the best results from these setups.
Table 1: Accuracies ($\uparrow$) on in-domain and out-of-domain datasets using Pythia-410M. Data size refers to the total number of unique training examples used. All experiments in this table share the same total training steps and learning rate schedule (see Section 5.2.1). Results surpassing full training are highlighted in bold. The ranges for confidence, perplexity, and learnability are chosen according to the best-performing intervals reported in prior research (Section 5.1). For one-shot selection methods (excluding S2L), we use representations from either step 1000 or the end of fine-tuning Pythia-410M on MathInstruct, and report the better results, while S2Lselects data based on training trajectories from a smaller Pythia-70M model.
Data In-domain Out-domain
Selection Size GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
(Pretrained) 2.0 1.6 10.1 4.6 2.3 2.5 1.4 3.3
Random 30K 3.3 6.2 15.0 8.2 15.0 15.1 1.6 9.4
50K 3.7 6.4 18.1 9.4 17.0 11.6 1.2 9.7
100K 5.9 7.6 22.0 11.8 20.5 20.8 2.7 13.3
Least Confidence 30K 2.7 1.3 18.0 7.0 13.7 3.3 1.4 6.7
50K 2.1 1.7 21.0 8.3 14.5 3.5 1.0 7.3
100K 2.5 3.3 23.5 9.8 20.8 6.3 3.7 10.0
Middle Perplexity 30K 5.3 3.7 16.2 8.4 14.2 8.7 1.2 8.2
50K 3.2 5.9 20.5 9.9 18.1 11.3 $\mathbf{5.1}$ 10.7
100K 5.4 7.2 20.9 11.2 $\mathbf{23.8}$ 15.3 3.3 12.6
High Learnability 30K 6.1 1.6 19.1 8.9 10.7 9.9 1.4 8.1
50K 6.1 2.1 18.6 8.9 14.5 14.0 2.1 8.9
Facility Locations 30K 4.2 7.7 10.0 7.3 11.8 13.8 1.2 8.1
50K 5.7 9.1 12.4 9.1 15.4 18.6 1.6 10.5
100K $\mathbf{7.4}$ $\mathbf{10.9}$ $\mathbf{30.5}$ $\mathbf{16.3}$ $\mathbf{26.2}$ $21.9$ $\mathbf{9.3}$ $\mathbf{17.7}$
DiverseEvol (Online) 30K 1.9 3.6 8.8 4.8 13.9 3.0 1.6 5.5
50K 1.6 4.2 12.0 5.9 10.6 7.3 1.9 6.3
100K 1.3 3.8 12.9 6.0 11.8 8.4 1.0 6.5
Confidence Curriculum (Online) 30K 4.2 6.3 15.4 8.6 18.9 16.5 1.4 10.4
50K 6.6 3.3 16.9 9.0 19.9 19.6 2.1 11.4
100K 4.6 6.3 17.1 9.3 21.0 15.2 1.8 11.0
S2L (Ours) 30K 5.0 $\mathbf{10.3}$ $\mathbf{29.4}$ $\mathbf{14.9}$ 20.6 17.3 7.8 15.1
50K $\mathbf{7.6}$ $\mathbf{9.7}$ $\mathbf{28.1}$ $\mathbf{15.1}$ 22.8 22.9 5.1 $\mathbf{16.0}$
100K $\mathbf{9.3}$ $\mathbf{11.8}$ $\mathbf{33.6}$ $\mathbf{18.3}$ $\mathbf{26.6}$ $\mathbf{26.5}$ $\mathbf{12.8}$ $\mathbf{20.1}$
None 262K 6.7 9.4 26.6 14.3 23.5 23.3 3.7 15.5
5.2 Mathematical Reasoning
Table 2: Accuracies ($\uparrow$) across model scales and model suites. Data selected by S2L with one small model (Pythia-70M) scale up to training larger models and can also generalize to other pretrained models. Results surpassing full training are highlighted in bold.
Target Model Fine-tuning Data GSM8K MATH NumGLUE Avg Diff to Full
Pythia-410M (Pretrained) 2.0 1.6 10.1 4.6 $-9.7$
S2L-50K(Ours) $\mathbf{7.6}$ $\mathbf{9.7}$ $\mathbf{28.1}$ $\mathbf{15.1}$ $+0.8$
Full-262K 6.7 9.4 26.6 14.3 -
Phi-2 (2.7B) (Pretrained) 53.4 16.1 34.9 $34.8$ $-20.3$
S2L-50K(Ours) 66.8 32.7 65.0 $54.8$ $-0.3$
Full-262K $68.3$ $32.6$ $64.3$ $55.1$
5.2.1 Training Settings
@@
For mathematical reasoning, we focus on fine-tuning using the MathInstruct dataset (Yue et al., 2023) with 262,040 training examples for its comprehensive coverage of diverse mathematical fields and its capability in training models to achieve state-of-the-art performance on the standard evaluation benchmarks (Section 5.2.2). We employ the open-source model suites Pythia (Biderman et al., 2023b) and a current state-of-the-art model Phi-2 (Li et al., 2023b) as our base models to validate our S2L algorithm across different scales and directly compare its performance against the state-of-the-art. Following the setup used in (Yue et al., 2023), we adopt a training regimen with a learning rate of 2e-5, a batch size of 128, a maximum length of 512, and a cosine scheduler with a 3% warm-up period. For all experiments on MathInstruct, we standardize the number of training steps to correspond to 3 epochs on the full dataset, also aligning with the setting in (Yue et al., 2023). This allows us to maintain a consistent training schedule across different methods and data budgets, ensuring fair and accurate comparisons of the quality of data selected by different algorithms.
Due to memory and computational constraints, for Facility Locations and DiverseEvol, we calculate pairwise similarity and perform greedy selection on a per-data-source basis. We found this per-source selection approach also yields benefits for S2L as different data sources within MathInstruct exhibit distinct common patterns in their training trajectories. Therefore, we implement S2L also on a per-source basis for MathInstruct, and recommend applying S2L per source when dealing with datasets composed of multiple data sources.
5.2.2 Evaluation Settings
@@
Datasets. We follow the framework established in (Yue et al., 2023) for a comprehensive assessment using several well-regarded datasets, including in-domain and out-of-domain datasets. For the in-domain datasets, we consider GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and NumGLUE (Mishra et al., 2022). For the out-of-domain datasets, we consider SVAMP (Patel et al., 2021), Mathematics (Davies et al., 2021), SimulEq (Koncel-Kedziorski et al., 2016). These datasets collectively span a diverse range of mathematical subjects, such as algebra, probability, number theory, calculus, and geometry. Additionally, some questions in these datasets require the application of commonsense, reading comprehension, and multi-step reasoning. All questions are open-formed.
@@
Metric. We use the standard evaluation metric for open-formed questions, exact match, which measures the model’s accuracy by comparing its generated answers against the correct solutions. For an answer to be considered correct, it must match the reference solution precisely.
5.2.3 Results
We present a comprehensive analysis of S2L in comparison to the full training baseline, data selection baselines, and the performance of the pretrained model (as a baseline to demonstrate the impact of fine-tuning).
Scaling the Data: SOTA algorithms fail with small data budgets while S2L stands out across data scales. In Table 1, we compare S2L against the baselines from Section 5.1 on Pythia-410M across varying data sizes. The training trajectories used by S2L are based on Pythia-70M, a model approximately 6x smaller than Pythia-410M. With the same number of training steps as the full training, S2L exceeds the full dataset’s performance using only 30K examples, only 11% of the full dataset. When the data budget is increased to 100K examples, S2L outperforms the full training by 4% in exact-match accuracy. It leads the runner-up baselines by an average of 4.7%, 4.6% and 2.4% with data budget 30K, 50K and 100K across all six evaluation datasets. We observe that, in this challenging fine-tuning setting with a domain shift, while state-of-the-art data selection algorithms like Facility Locations (Bhatt et al., 2024) and High Learnability (Zhou et al., 2023b) have decent performance with a large enough data budget (i.e., 100K), all SOTA algorithms except S2L cannot even outperform the random sampling baseline when the allowed data size is small (i.e., 30K). Unlike the existing algorithms, S2L consistently outperforms all baselines and even full training across all data sizes. Note that compared to the runner-up algorithm in this setting, Facility Locations, the cost of S2L is much lower in both training the reference model and data selection stages (Figure 4), and therefore more scalable to both larger target models or larger data sizes.
@@ 5.2.3 Results
Scaling the Model: Data selected by S2L can transfer to larger models in different model suites. We also test whether this subset, chosen using Pythia-70M, can effectively train larger models beyond 410M and models outside the Pythia suite. Since larger models are easier to overfit, we randomly sample 50 examples from each data source as the validation set and use the best model with the lowest validation loss evaluated every 1000 training iterations. Our experiments with Phi-2 reveal that fine-tuning on only 50K S2L-selected data again outperforms full dataset training on the most challenging MATH (Hendrycks et al., 2021) benchmark improving the pretrained Phi-2 by $16.6\%$ and is more data efficient than training on the full MathInstruct dataset to get the same performance.
(a) Examples in the cluster shown in Figure 5(a): open-formed algebra.
(b) Examples in the cluster shown in Figure 5(b): reading comprehension + coding.
(c) Examples in the cluster shown in Figure 5(c): multiple-choice + multi-step reasoning.
Figure 5: Examples (MathInstruct) with similar training trajectories have the same question format. Intuitively, training on one of the questions helps the model learn how to solve the other questions with the same format, and therefore their losses have similar trends. Sampling from all clusters and prioritizing the small ones help ensure different types of examples can be learned. See the full version in Appendix B.
@@
Examples in Clusters Encode the Same Knowledge/Skill. In Figure 5, we compare actual training examples in MathInstruct that get clustered together due to their similar training trajectories on the small Pythia-70M model. We observe that examples in the same cluster are of the same type and related to the same knowledge/skill: the cluster shown in Figure 4(a) is comprised of open-formed algebra questions, examples in Figure 4(b) requires extracting useful information from long text and write programs to answer the question, and Figure 4(c) shows a cluster of multiple choice questions that require multi-step reasoning. Therefore, by sampling from different clusters, we make sure the selected examples cover the knowledge required for all topics and skills required for all types of questions.
5.3 Clinical Text Summarization
@@
S2L can improve data efficiency not only for fine-tuning data in the mathematical domain but also for other specialized domains. In this subsection, we focus on the task of clinical text summarization within the context of radiology reports. This task involves processing the detailed analysis and results listed in the findings section of a radiology report and distilling them into a concise impression section. Such summaries are crucial for providing clinicians with quick and actionable insights from radiological studies.
Table 3: A synthetic radiology report (MRI of the brain), generated by the GPT-4 model (Achiam et al., 2023) to demonstrate the typical data format and content used in the clinical text summarization task. It is not suitable for clinical or diagnostic use.
Findings
The brain parenchyma demonstrates normal morphology with no evidence of mass effect or midline shift. No acute infarcts are seen on diffusion-weighted images. There are no signs of intracranial hemorrhage. Mild generalized cerebral atrophy is noted. The ventricles and sulci appear within normal limits for the patient’s age. The pituitary gland and sella turcica are unremarkable. There are no abnormal signal intensities within the brain parenchyma. The orbits, paranasal sinuses, and mastoid air cells are clear.
Impression
Normal MRI of the brain. Mild cerebral atrophy, likely age-related. No acute intracranial pathology.
@@
Dataset & Setup. We use the MIMIC-III dataset (Johnson et al., 2016), a comprehensive collection of radiology reports and findings authored by attending physicians in routine clinical care. We use the same preprocessing procedures as (Delbrouck et al., 2023; Demner-fushman et al., 2023) to extract the findings and impression sections and remove invalid reports. Given that access to MIMIC-III requires specific credentials, we provide a synthetic example of a radiology report generated by GPT-4 (Achiam et al., 2023) for illustrative purposes in Table 3. We employ the Pythia-1B model and keep the training setting consistent with the mathematical reasoning task.
Evaluation. Our evaluation of the generated clinical summaries on the test split of the MIMIC-III dataset uses three key metrics, as suggested in (Van Veen et al., 2023; Tu et al., 2023). (1) BLEU (Papineni et al., 2002) calculates the overlap between the generated and reference texts using short sequences of words. (2) ROUGE-L (Lin, 2004) examines the longest sequence of words that appears in both texts. (3) BERTScore (Zhang et al., 2020), utilizing BERT’s contextual embeddings, evaluates how semantically similar the generated text is to the reference. These metrics together offer a comprehensive evaluation, from basic word-level accuracy to deeper semantic alignment, ensuring our summaries are not only precise in language but also meaningful and coherent in the context of clinical information. We compare S2L to random selection, a surprisingly strong baseline as evidenced in Table 1, to check the validity of the data selection problem on this dataset and then compare it to training on the full dataset to assess its effectiveness.

### v2

@@ 5 Experiments
In this section, we present the comprehensive experiments conducted to evaluate the efficacy of the proposed data selection method, SmallToLarge (S2L), across two challenging domains (mathematical reasoning and clinical text summarization).
@@ 5.1 Baselines
We systematically compare S2L against a comprehensive set of open-sourced data selection methods. These methods are categorized based on the type of representation they use and selected as the most representative or best-performing methods as identified in prior work. These include: (1) Random Sampling; selecting examples with the (2) Least Confidence [4] or (3)Middle Perplexity [34]; (4) High Learnability, determined by the loss decrease before and after full fine-tuning [68]; and (5) Facility Locations selection based on hidden states [4]. Additionally, we incorporate one online selection techniques: (6) Confidence Curriculum proposed by Varshney et al., which selects examples with decreasing confidence during the training. Given that the optimal reference model may vary for each one-shot selection method, we ensure a fair comparison by adopting the approach used in [34], which runs each method with both the fully fine-tuned target model and an early fine-tuning checkpoint as the reference model. We report the best results from these setups.
5.2 Specialized Domain 1: Mathematical Reasoning
@@
Training Settings. We focus on fine-tuning using the MathInstruct dataset [64] with 262,040 training examples for its comprehensive coverage of diverse mathematical fields and its capability in training models to achieve state-of-the-art performance on the standard evaluation benchmarks. We employ the open-source model suites Pythia [6], Phi-2 [28], Llama-2 [50] as our base models to validate our S2L algorithm and directly compare its performance against the state-of-the-art.
@@
Evaluation Datasets. We follow the framework established in [64] for a comprehensive assessment using several well-regarded datasets, including in-domain and out-of-domain datasets. For the in-domain datasets, we consider GSM8K [11], MATH [19], and NumGLUE [36]. For the out-of-domain datasets, we consider SVAMP [38], Mathematics [13], SimulEq [25]. These datasets collectively span a diverse range of mathematical subjects, such as algebra, probability, number theory, calculus, and geometry. Additionally, some questions in these datasets require the application of commonsense, reading comprehension, and multi-step reasoning. All questions are open-formed.
@@
Evaluation Metric. We use the standard evaluation metric for open-formed questions, exact match, which measures the model’s accuracy by comparing its generated answers against the correct solutions. For an answer to be considered correct, it must match the reference solution precisely.
More details about the settings and baseline implementations can be found in Appendix B.
(a) In-domain Avg
(b) Avg
Figure 4: Data Scaling: Accuracies ($\uparrow$) on in-domain and out-of-domain datasets using Pythia-410M. Data size refers to the total number of unique training examples used. All experiments in this table share the same total training steps and learning rate schedule (see Section 5.2). See breakdowns in Figure 14.
Table 1: Less Data, Same Compute: Zero-shot accuracies (%, $\uparrow$) when S2L and the baselines select 50K data to train with the same number of iterations as the full-data training. Results surpassing full training are highlighted in bold. Figure 4 follows the same setting but uses the Pythia-410M model.
Target Fine-tuning In-domain Out-domain
Model Data GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
Phi-2 (2.7B) (Pretrained) 53.4 16.1 34.9 $34.8$ 67.9 31.1 27.4 38.5
Random 67.9 30.1 60.7 52.9 77.1 51.2 37.5 54.1
High Learnability 59.4 25.2 62.1 48.9 76.6 41.8 27.2 48.7
Middle Perplexity 66.4 29.5 54.1 50.0 74.8 50.4 39.8 52.5
Least confidence 61.7 24.7 67.0 51.1 76.5 43.3 52.5 54.3
Facility Locations 66.2 31.3 62.4 53.3 74.4 58.4 34.6 54.5
S2L(Ours) $\mathbf{69.1}$ $\mathbf{32.6}$ $\mathbf{65.7}$ $\mathbf{55.8}$ $\mathbf{79.6}$ 56.4 40.1 57.3
Full-262K $68.3$ $32.6$ $64.3$ $55.1$ 78.4 58.4 44.2 $\mathbf{57.7}$
@@
5.2.1 Setting 1: Less Data for Better Models
In the first setting, we standardize the number of training steps to correspond to 3 epochs on the full dataset, aligning with [64]. This allows us to maintain a consistent training schedule across different methods and data budgets, ensuring fair and accurate comparisons of the quality of data.
Scaling the Data: SOTA algorithms fail with small data budgets while S2L stands out across data scales. In Figure 4, we compare S2L against the baselines from Section 5.1 on Pythia-410M across varying data sizes. The training trajectories used by S2L are based on Pythia-70M, a model approximately 6x smaller than Pythia-410M. With the same number of training steps as the full training, S2L exceeds the full dataset’s performance using only 30K examples, only 11% of the full dataset. It leads the runner-up baselines by an average of 4.7%, 4.6% and 2.4% with data budget 30K, 50K and 100K across all six evaluation datasets. While state-of-the-art data selection algorithms like Facility Locations [4] and High Learnability [68] have decent performance with a large enough data budget (i.e., 100K), all SOTA algorithms except S2L cannot even outperform the random sampling baseline when the allowed data size is small (i.e., 30K). Unlike the existing algorithms, S2L consistently outperforms all baselines and even full training across all data sizes. Note that compared to the runner-up algorithm in this setting, Facility Locations, the cost of S2L is much lower in both training the reference model and data selection stages (Figure 5), and therefore more scalable to both larger target models or larger data sizes.
@@ 5.2.1 Setting 1: Less Data for Better Models
Scaling the Model: Data selected by S2L can transfer to larger models in different model suites. We also test whether this subset, chosen using Pythia-70M, can effectively train larger models beyond 410M and models outside the Pythia suite. As shown in Table 1, our experiments with Phi-2 reveal that fine-tuning on only 50K S2L-selected data again outperforms full dataset training on the most challenging MATH [19] benchmark improving the pretrained Phi-2 by $16.6\%$ and is more data efficient than training on the full MathInstruct dataset to get the same performance.
Figure 6: Distribution of the coverage of top-1 topic in each cluster. Taller bars on the right end of the plot indicate clusters with a higher concentration of a single topic and therefore suggest a grouping of similar examples.
5.2.2 Setting 2: Less Data for Faster Training
The second setting we consider is when fixing the number of times each example can be seen over the entire course of training, directly translating smaller datasets into reduced training and storage costs. This is particularly beneficial for large models that would require extensive training times without data selection. By experimenting with models of larger sizes than the previous setting, we observe in Table 2 that S2L can achieve comparable performance to full-data training when using only 50% data and thereby reducing both the data storage space and the training time by half.
Table 2: Less Data, Same Epochs: Zero-shot accuracies (%, $\uparrow$) when S2L trains 50% data for the same number of epochs as the full-data training. S2L can achieve comparable performance to full-data training while reducing both the data storage space and the training time by half.
Target Fine-tuning In-domain Out-domain
Model Data GSM8K MATH NumGLUE Avg SVAMP Mathematics SimulEq Avg
Phi-3-mini (3.8B) (Pretrained) 74.5 26.5 52.1 51.1 83.7 44.3 34.8 52.7
S2L-50%(Ours) 76.3 42.5 76.4 65.1 83.8 62.1 51.6 65.4
Full 76.4 42.9 75.3 64.9 84.6 60.2 51.9 65.2
Llama-2-7B (Pretrained) 3.1 4.2 16.5 7.9 14.1 8.3 2.3 8.1
S2L-50%(Ours) 53.3 28.9 65.0 49.1 65.1 45.2 31.9 48.2
Full-262K [64] 52.2 30.4 60.5 47.7 65.3 43.9 50.2 50.4
5.2.3 Why is S2L So Effective?
@@
Examples in Clusters Encode the Same Knowledge/Skill. In Appendix C, we compare actual training examples in MathInstruct that get clustered together due to their similar training trajectories on the small Pythia-70M model. We observe that examples in the same cluster are of the same type and related to the same knowledge/skill, e.g., open-formed algebra questions (Figure 15), examples requiring extracting useful information from long text and writing programs (Figure 16), and multiple choice questions that require multi-step reasoning (Figure 17), etc. Therefore, by sampling from different clusters, we make sure the selected examples cover the knowledge required for all topics and skills required for all types of questions.
Loss Trajectories can Capture the Similarity Between Data Points As Much As Embeddings of a Fully Fine-tuned Model. We conducted a quantitative analysis to assess how effectively S2L identifies similar examples using loss trajectories from a small model. Assuming math problems under the same topic require similar knowledge and share question formats, we used unknown topic labels during S2L’s data selection to check if each cluster predominantly contains a single topic. By calculating the fraction of the most common topic in each cluster and plotting this in Figure 6 (with K=100, excluding clusters of size one), we compared the loss trajectory clusters from S2L (in blue) against those from the embeddings of a fully fine-tuned Phi-2 model (in orange)—considered the ground truth for similarity. Results show that most clusters formed by S2L using the Pythia-70M model are based on a single topic and capture topic similarities more effectively than those from the Phi-2 model’s embeddings. This analysis not only confirms the homogeneity within S2L clusters but also highlights the computational efficiency of using loss trajectories on small models to identify representative examples.
5.3 Specialized Domain 2: Clinical Text Summarization
@@
S2L can improve data efficiency not only for fine-tuning data not only in mathematics but also in other specialized domains. This subsection explores its application to clinical text summarization within radiology reports. This task involves processing the detailed analysis and results listed in the findings section of a radiology report and distilling them into a concise impression section. Such summaries are crucial for providing clinicians with quick and actionable insights from radiological studies.
@@
Dataset & Setup. We use the MIMIC-III dataset [21], a comprehensive collection of radiology reports and findings authored by attending physicians in routine clinical care. We use the same preprocessing procedures as [14, 15] to extract the findings and impression sections and remove invalid reports. Given that access to MIMIC-III requires specific credentials, we provide a synthetic example of a radiology report generated by GPT-4 [2] for illustrative purposes in Table 3. We employ the Pythia-1B model and keep the training setting consistent with the mathematical reasoning task.
@@
Evaluation. Our evaluation of generated clinical summaries on the MIMIC-III dataset’s test split employs three key metrics as recommended in [52, 51]: (1) BLEU [37], which measures word sequence overlap between the generated and reference texts; (2) ROUGE-L [29], assessing the longest common word sequence; and (3) BERTScore [65], evaluating semantic similarity using BERT’s contextual embeddings. These metrics together offer a comprehensive evaluation, ensuring our summaries are not only precise in language but also meaningful and coherent in the context of clinical information. We compare S2L to random selection, a surprisingly strong baseline as evidenced in Section 5.2, to check the validity of the data selection problem on this dataset and then compare it to training on the full dataset to assess its effectiveness.